OpenStudy (anonymous):

Me. Cooper invested a sum of money at 6%. He invested a second sum, $150 more than the first 3%. If his total annual income was $54, how much did he invest at 3%?

OpenStudy (amistre64):

a = b-e ax +by = i (b-e)x + by = i bx -ex +by = i b(x+y) = i + ex sum @ 3% = (54 + 150(.06))/(.06 + .03) if i see it right, have to dbl chk

OpenStudy (amistre64):

700@ 3% is what I get 700(.03) + 550(.06) = 54 21 + 33 = 54 54 = 54

OpenStudy (anonymous):

Kk ty